OpenStudy (anonymous):

A line passes through the points (-10, -4) and (-1,2). What is the slope of the line that passes through the two points? What is the y-intercept?

OpenStudy (johnweldon1993):

\[m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] You have 2 points (-10, -4) and (-1,2) (x1 , y1) (x2,y2) plug those in and solve for the slope

OpenStudy (anonymous):

we can solve this by using the slope formula: \[m=\frac{ y2-y1 }{ x2-x1 }\] this gives us our slope to find the y-intercept, we should first find the equation for the line with point slope form: \[y-y1=m(x-x1)\] once you substitute terms, just solve for y.